### v1

# Abdellah El Mekkiλ          Houdaifa Atouδ          Omer Nacarψ Shady Shehataξ,γ          Muhammad Abdul-Mageedλ,ξ,γ

Abstract
Enhancing the linguistic capabilities of Large Language Models (LLMs) to include low-resource languages is a critical research area. Current research directions predominantly rely on synthetic data generated by translating English corpora, which, while demonstrating promising linguistic understanding and translation abilities, often results in models aligned with source language culture. These models frequently fail to represent the cultural heritage and values of local communities. This work proposes a methodology to create both synthetic and retrieval-based pre-training data tailored to a specific community, considering its (i) language, (ii) cultural heritage, and (iii) cultural values. We demonstrate our methodology using Egyptian and Moroccan dialects as testbeds, chosen for their linguistic and cultural richness and current underrepresentation in LLMs. As a proof-of-concept, we develop NileChat, a 3B parameter LLM adapted for Egyptian and Moroccan communities, incorporating their language, cultural heritage, and values. Our results on various understanding, translation, and cultural and values alignment benchmarks show that NileChat outperforms existing Arabic-aware LLMs of similar size and performs on par with larger models. We share our methods, data, and models with the community to promote the inclusion and coverage of more diverse communities in LLM development. [^1] https://github.com/UBC-NLP/nilechat.
[Uncaptioned image] NileChat: Towards Linguistically Diverse and Culturally Aware LLMs for Local Communities
Abdellah El Mekkiλ          Houdaifa Atouδ          Omer Nacarψ Shady Shehataξ,γ          Muhammad Abdul-Mageedλ,ξ,γ λThe University of British Columbia           δUM6P           ψPSU ξMBZUAI           γInvertible AI {abdellah.elmekki,muhammad.mageed}@ubc.ca
{strip}[Uncaptioned image]
Figure 1: Our proposed framework for data augmentation tailored to low-resource local communities. (a) Workflow for generating educational data in the target language through machine translation. (b) Workflow for generating diverse text genres in the target language by simulating scenarios that incorporate local cultural contexts and persona descriptions. (c) Retrieval process for augmenting local cultural knowledge by parsing web pages representing specific local cultural concepts through web search.
1 Introduction
Large Language Models (LLMs) have advanced rapidly, enabling remarkable proficiency across many tasks. Yet, this success is unevenly distributed across languages, with substantial performance disparities observed for non-English languages, particularly low-resource languages and dialectal variants Navigli et al. (2023). A primary factor underlying this discrepancy is the limited representation of diverse multilingual data within the foundational pre-training corpora of these models which favors high-resource languages spoken in regions with high economic influence Bender (2011); Joshi et al. (2020).
Beyond linguistic limitations, a more profound challenge is the inherent risk of cultural encapsulation Wrenn (1962) in LLMs. Cultural encapsulation refers to an unconscious tendency to operate within one’s own cultural lens, leading to misunderstanding or avoidance of differing perspectives and values. As LLMs are optimized to replicate patterns in their training data—predominantly sourced from specific cultural contexts (e.g., Western, English-speaking)—they risk internalizing and propagating these dominant perspectives as the norm Dwivedi et al. (2023); Tao et al. (2024); Wang et al. (2024); Naous et al. (2024). The significance of cultural context cannot be overstated. As Edward Sapir noted:
"No two languages are ever sufficiently similar to be considered as representing the same social reality. The worlds in which different societies live are distinct worlds, not merely the same world with different labels attached." - Sapir (1929)
This cultural bias is compounded by a fundamental mismatch: LLMs typically process data through a language-centric lens, whereas human communities are structured around shared social ties, perspectives, and values MacQueen et al. (2001). Current LLMs adaptation techniques for new languages or communities Gurgurov et al. (2024); Joshi et al. (2025) often fall short in bridging this cultural divide, especially for low-resource communities Naous et al. (2024). For instance, machine translation, while useful for generating synthetic data to boost linguistic coverage Joshi et al. (2025); Shang et al. (2025); Wang et al. (2025), primarily addresses the linguistic deficit. The translated content often retains the source language’s cultural perspective, failing to incorporate authentic local nuances crucial for genuine interaction. Supervised fine-tuning (SFT) on target language data Gala et al. (2024); Shang et al. (2025) can align models to specific tasks, but small datasets may not reshape deep-seated cultural biases from pre-training Rystrøm et al. (2025) and can encourage hallucination with new factual data Gekhman et al. (2024). While continued pre-training with culturally rich data could mitigate these issues, it faces a critical bottleneck for low-resource contexts: the scarcity of such high-quality digital texts.
This paper addresses the critical need to adapt multilingual LLMs to low-resource language communities by jointly considering their linguistic characteristics and cultural heritage & values. We propose a novel pipeline (illustrated in Figure 1) focused on data augmentation for continued pre-training. Our approach combines controlled synthetic data generation (Section 3.1.2) with retrieval (Section 3.1.3) methods. To address linguistic adaptation, we translate English pre-training data into the target local language focusing only on high-quality data from the educational domain (Section 3.1.1). Crucially, to imbue cultural relevance, we generate diverse texts reflecting specific cultural heritage concepts (e.g., food, celebrations, proverbs) using local persona descriptions (Section 3.1.2) reflecting the local cultural values. We demonstrate our method on the Moroccan and Egyptian Arabic dialects as low-resource testbeds. We further pre-train a multilingual LLM on a curated mix of real and synthetic data, evaluating its performance on tasks involving language understanding, translation, and alignment with cultural knowledge and values. Our findings show that the adapted model significantly outperforms baseline and existing models that are even bigger in size on most evaluation tasks.
The main contributions of this work are: (i) A novel framework for augmenting pre-training corpora tailored to local communities. This framework considers their unique linguistic features, cultural heritage, and values by leveraging a teacher LLM. (ii) The public release of new datasets, representing the largest publicly available corpora for Egyptian and Moroccan Arabic dialects. These resources are intended to foster further research in these under-resourced languages. (iii) The development and public release of NileChat, a robust 3-billion parameter LLM. This model demonstrates proficiency in both Egyptian and Moroccan dialectal Arabic (using Arabic script and Arabizi) while maintaining strong performance in Modern Standard Arabic, French, and English.
2 Related Work
Adaptation of LLMs.
LLMs, despite general strengths, often require adaptation for specific languages, domains, or cultures Bang et al. (2023); AlKhamissi et al. (2024); Naous et al. (2024); Song et al. (2025). Adaptation techniques include prompt engineering Shen et al. (2024), SFT on culturally specific datasets Huang et al. (2024), and continued pre-training on target-specific data Fujii et al. (2024); Huang et al. (2024). A key challenge, especially for SFT-based cultural adaptation, is the scarcity of comprehensive cultural datasets, hindering alignment with under-represented communities Ahmad et al. (2024); Shen et al. (2024).
Synthetic Data Augmentation for LLMs.
To address data limitations, synthetic data augmentation has shown promise in improving LLM performance Ge et al. (2024); Li et al. (2024); Joshi et al. (2025). Machine-translated data, for instance, can enhance capabilities in new languages Joshi et al. (2025); Shang et al. (2025), and persona-driven synthetic data generation has also yielded performance gains Ge et al. (2024) and aided in tasks like assessing LLM political alignment Bernardelle et al. (2024). However, synthetic data can sometimes degrade performance Seddik et al. (2024), necessitating best practices for its use Liu et al. (2024).
Arabic LLMs.
In Arabic LLM development, models are either trained from scratch Billah Nagoudi et al. (2023); Sengupta et al. (2023) or adapted from existing ones Huang et al. (2024); Bari et al. (2025); Team et al. (2025a). A common method involves translating English data to Arabic, which, however, can introduce cultural biases from the source language Sengupta et al. (2023); Naous et al. (2024). Recent work on dialectal Arabic, such as translating instructions into Moroccan dialect for SFT, has improved generation tasks Shang et al. (2025). Yet, enhanced performance on standard tasks does not guarantee cultural awareness. While models like AceGPT Huang et al. (2024) and Fanar Team et al. (2025a) aim for cultural cognizance, our work uniquely focuses on adapting existing LLMs to a local community by deeply integrating its specific linguistic features, cultural heritage, and values, building upon these prior advancements.
3 Methodology
In this work, we investigate the potential of pre-training data to imbue LLMs with the specific local characteristics of under-represented communities. We conceptualize these characteristics along three primary dimensions Geertz (1977); Anderson (1991); Bourdieu and Thompson (1991); Higgins and Douglas (2020); Stanlaw and Adachi (2025): (i) Language: Encompassing dialectal nuances, idiomatic expressions, and linguistic structures unique to the community. (ii) Cultural Heritage: Reflecting the customs, traditions, social norms, historical context, and common knowledge prevalent within the community. (iii) Cultural Values: Capturing the ethical standpoints, belief systems, and societal priorities that define the community. We refer to these three dimensions as Language-Heritage-Values dimensions, LHV for short. While we do not posit these as exhaustive of the attributes of a given community, we employ them as a vehicle to approximate the LLM communication and information needs at local levels. To ground our investigation, we focus on two low-resource varieties of Arabic—The Egyptian Arabic (EGY) and Moroccan Arabic (MOR)—. These dialects serve as our primary case studies for evaluating the methods proposed herein.
3.1 Data Augmentation
The construction of LLMs that can serve a specific population fundamentally depends on the availability of representative data. Recognizing the acute scarcity of publicly available pre-training corpora for many low-resource languages, including EGY and MOR, we propose a novel data production method encapsulating the LHV dimensions of a given country-level population. As depicted in Figure 1, our approach leverages three complementary strategies intended to collectively capture the LHV dimensions: (a) machine translation (MT), (b) controlled-generation and (c) retrieval. We explain these next.
3.1.1 MT for Knowledge and Fluency
To ensure linguistic fluency and coherence, we translate structured educational content from English into the target low-resource language using a specialized teacher model. Our pipeline preserves original formatting and includes filtering to remove unreliable translations identified by repetitive n-grams. We use educational materials for their topical breadth (including subjects such as education, history, health, medicine, and biology).
3.1.2 Controlled Generation for Cultural Heritage and Cultural Values
Linguistic fluency, while a foundational capability for LLMs, does not inherently guarantee their awareness of, or alignment with, the culture and values of a specific target community Naous et al. (2024). To bridge this gap, we employ controlled generation. For controlled generation, we use the teacher LLM to generate diverse texts in the target language. These texts are specifically designed to discuss local topics, which are identified from articles sourced from local news websites or the target country’s Wikipedia portal. Furthermore, the generated content is crafted to reflect distinct personas, each defined by a profile encompassing specific moral values, demographic characteristics, and socioeconomic attributes. Our approach integrates four key components to achieve this:
Local Contextual Information. We ground our synthetic data by incorporating local context drawn from news websites within the target communities. These sources provide relevant contextual information and do not necessarily need to be in the target local language.
Core Cultural Heritage Concepts. We integrate key regional cultural elements, such as cuisine, landmarks, and celebrations, by extracting relevant articles from country-specific Wikipedia portals.
Linguistic and Cultural Expressions. To authentically capture local idiomatic styles, we collect common expressions, proverbs, idioms, dialogues from TV programs, and local terminology, pairing each with English translations for accuracy.
Representative Personas. We develop representative personas reflecting local moral, demographic, and socioeconomic attributes by leveraging data from the World Values Survey (WVS) EVS/WVS (2024). Selected survey responses are transformed into textual descriptions, which are further refined by an LLM to create concise and coherent persona profiles (see Figure 2).
To produce diverse text genres for pre-training, we combine data points from the four listed components into a unified prompt to guide the teacher LLM. This prompt instructs the LLM to generate varied text outputs in the target low-resource language, explicitly integrating the selected persona’s values, the specified cultural concepts, and provided linguistic cues. Specifically, we focus on generating the following genres: stories, personal essays, blog posts, reviews, and conversations. An example of this process is depicted in Figure 1 (b).
3.1.3 Retrieval for Local Cultural Heritage
This method involves querying a search engine using a pre-defined list of cultural concepts that span multiple cultural categories. For each concept, we extract the top $20$ search results, systematically excluding social media platforms. The textual content from the retrieved web pages is then parsed and extracted using Trafilatura Barbaresi (2021).
Figure 2: Pipeline for generation of persona descriptions using the WVS.
3.2 Model Training
We evaluate our proposed method on Egyptian (EGY) and Moroccan (MOR) Arabic dialects. Despite their large speaker populations, these dialects remain low-resource, underscoring the need for specialized language models. We select Command R+ Cohere Labs (2024) (104B) as our teacher model, as it demonstrates reasonable text-generation capabilities in both target dialects. Additionally, Command R+ provides open weights, enabling us to efficiently generate or translate extensive datasets without incurring API costs.
3.2.1 Continued Pre-Training
Data.
We generate pre-training data for EGY and MOR using the methods outlined in Section 3.1. Our approach involves three main components: (i) MT Data. We employ our teacher model to translate English educational content into both dialects. Specifically, we sample 5.5 million texts from the Fineweb-edu dataset Penedo et al. (2024) and translate them into EGY and MOR. (ii) Controlled Generation Data. We craft tailored prompts incorporating personas, local cultural contexts, dialectal glosses, expressions, and utterances to instruct the LLM in generating diverse genres of text. For persona descriptions, we generate 1,200 descriptions based on data from Egyptian and Moroccan participants in the WVS. For local news context, we leverage an in-house corpus comprising approximately 1.5 million Egyptian and 800,000 Moroccan news articles, originally published in MSA by local news websites. Additionally, we include 25,000 Egyptian and 49,000 Moroccan Wikipedia articles. For dialectal glosses, expressions, and utterances, we draw from publicly available resources on EGY and MOR proverbs and idiomatic expressions, each accompanied by English explanations. We further augment this with an in-house dataset of 600 dialectal utterances from Egyptian and Moroccan television shows paired with English translations, as well as 4,000 dialect-to-English word pairs for each dialect from the Gatitos dictionary Jones et al. (2023). (iii) Retrieval Data. For information retrieval, we query the Brave Search API [^2] https://brave.com/search/api/ using 6,500 cultural concepts from Morocco and 4,500 cultural concepts from Egypt. These concepts represent the ten cultural heritage categories in the set {food, clothes, landmarks, festivals & celebrations, geography, handicrafts, architecture, auna, flora, music}.
The generated dataset comprises approximately 5.5 million educational articles for both EGY and MOR. Additionally, for EGY, it includes approximately 300,000 samples for each category of conversions, personal essays, blog posts, reviews, and stories. For MOR, there are approximately 150,000 samples for each of these same categories. These latter categories represent the LHV dimensions (§3.1). Table A.3 presents a sample of these texts.
A filtering process using a repetitive n-gram filter removed 3.97% of the data. We also conducted a dialectness check on the generated data using ALDi Keleg et al. (2023). The average dialectness scores for the EGY and MOR educational articles are $0.45$ and $0.32$, respectively. In contrast, for the texts focused on cultural heritage and values, the average dialectness scores are higher, at $0.84$ for EGY and $0.72$ for MOR. We attribute the lower dialectness levels in the educational articles to the prevalence of scientific terms that often lack direct equivalents in EGY and MOR, and were therefore retained in MSA. We convert 1.5M and 0.5M from this generated data to Arabizi for both EGY and MOR, respectively. For retrieval, we collect 110,000 and 30,000 articles about cultural heritage to both EGY and MOR.
Our final pre-training dataset is a mixture of our generated and retrieved data, combined with pre-existing publicly available data for these dialects, MSA, English, French, Math, and Code. Our objective is to preserve the data distribution of the base model’s pre-training data to mitigate catastrophic forgetting Luo et al. (2025). The resulting pre-training dataset comprises 98.57 billion words, and its composition is detailed in Table A.1.
Compute.
We used a cluster of 4$\times$×A100 80GB GPUs for 1,096 hours to create our augmented pre-training dataset using the listed inputs.
Continued Pre-training.
Rather than pretraining an LLM from scratch, we continue pretraining Qwen-2.5-3B Qwen et al. (2025) with our data. We select this model due to its competitive performance and good tokenizer compression ratio on MSA. We continue pretraining the full model (3.1B parameters) for one whole epoch, which took 750 hours on 4$\times$×A100 80GB GPUs. More details about the base model selection and the training are in Appendix B.1.
3.2.2 Supervised Fine-Tuning
Data.
To adapt our pre-trained model for instruction following, we perform supervised fine-tuning (SFT). Due to the scarcity of SFT datasets for EGY and MOR, we construct a comprehensive training set. This process involves several key steps: (i) translation of SmolTalk dataset Allal et al. (2025) into MOR, EGY, French, and MSA using the teacher LLM; (ii) generation of synthetic dialectal question-answer pairs using our the retrieved dataset of local Egyptian and Moroccan cultural heritage; [^3] This data is initially created in Arabic script, and a portion is subsequently converted to Arabizi. (iii) incorporation of the Darija-SFT-Mixture MOR dataset provided by Shang et al. (2025); and (iv) translation of TULU-V2-mix dataset Ivison et al. (2023) into EGY. Finally, (v) this consolidated SFT dataset is augmented by converting understanding and generation tasks from the training sets of the ORCA Elmadany et al. (2023) and Dolphin Nagoudi et al. (2023) benchmarks into instruction-response formats. The final composition of our instruction dataset is in Table A.2.
Fine-Tuning.
For model fine-tuning, we follow recent approaches Ramé et al. (2024); Dang et al. (2024) that leverage model merging techniques to produce models effective across multiple languages or tailored for particular tasks. Specifically, we fully fine-tuned two separate variants of the base model—one specialized for MOR and the other for EGY—each trained on its respective dialectal data in both Arabic script and Arabizi (plus an amount of shared data between the two variants from the other languages; see B.2). We fine-tune each dialect-specific model for two epochs and employ weighted linear averaging Aakanksha et al. (2024) for merging, dubbing our merged model NileChat.
More information about our model merging is in B.2 and the prompts used for generating and translating our pre- and fine-tuning datasets is in E.
4 Experiments
4.1 Evaluation Tasks
We employ a comprehensive evaluation framework to measure the performance of NileChat for EGY and MOR. This framework enables comparison with our baseline and other LLMs across multiple capability dimensions: Understanding, cultural knowledge, translation, and value alignment.
Understanding.
We evaluate understanding capabilities using MMLU Hendrycks et al. (2021), HellaSwag Zellers et al. (2019), and Belebele Bandarkar et al. (2024) benchmarks, each adapted to both EGY and MOR dialects. For MOR, we directly employ the MMLU and HellaSwag versions provided by Shang et al. (2025). For EGY, we follow the translation pipeline described in Shang et al. (2025), translating the English and MSA MMLU tasks and the English HellaSwag dataset into EGY using our teacher model. For the Belebele benchmark, we utilize the official Moroccan and Egyptian dialect sets. Evaluations are conducted in both zero-shot and 3-shot scenarios, using accuracy as our performance metric.
Cultural Knowledge.
To assess cultural knowledge specific to Morocco and Egypt, we utilize the publicly available test set from the Palm benchmark Alwajih et al. (2025), focusing on these two countries. We adopt an LLM-as-Judge methodology Zheng et al. (2023), employing Gemma-3-27b Team et al. (2025b) to rate the correctness of model-generated responses compared to ground-truth answers on a scale from 0 to 10. The final evaluation score is calculated as the average correctness across all responses.
Translation.
[…]
Value Alignment.
To assess alignment with societal values, we adapt WVS questions into a multiple-choice format (expressed in the local language). The questions are categorized into 13 dimensions such as Economic Values (EcoV), Ethical Values (EthV), and Happiness and Wellbeing (HW). [^4] See Appendix C.1 for the full list. We use the Social Value Alignment (SVA) metric (Lee et al., 2024), which measures alignment using the distribution of survey responses. A model’s alignment score for each question corresponds to the proportion of participants who chose the model-predicted option, averaged across all questions for the final score.
Baseline Models.
We compare NileChat against a set of $17$ instruction-tuned LLMs known for their strong capabilities in Arabic, capped at 13B parameters (see full list in Table C.1 and Appendix C.2 for details). [^5] We also evaluate our translation performance against an NLLB-200’s 3.3B variant Team et al. (2022).
4.2 Results and Discussion
Model MMLU HellaSwag Belebele Palm
EGY MOR EGY MOR EGY MOR EGY MOR
Less than 7B Qwen3-1.7B 28.53 28.53 28.44 27.47 22.89 22.89 3.61 2.12
ar-stablelm-2-chat 41.56 40.36 34.79 33.45 38.89 36.11 4.20 3.62
Atlas-Chat-2B 42.61 44.87 29.66 34.74 50.56 55.67 3.16 3.42
Llama-3.2-3B-Instruct 40.68 37.54 29.16 28.27 45.44 35.89 3.21 2.28
gemma-3-4b-it 40.79 32.70 34.21 31.35 37.33 34.22 7.61 5.42
Qwen3-4B 28.61 28.54 30.28 29.04 22.89 22.89 4.51 2.71
Qwen2.5-3B-Instruct 43.37 44.43 31.62 29.58 51.33 41.44 2.86 2.31
NileChat (3B) 57.56 57.36 37.97 39.33 72.67 70.33 5.72 5.86
More than 7B AceGPT-7B-chat 40.29 37.57 33.27 30.47 32.67 32.00 5.58 3.93
ALLaM-7B-Instruct 60.04 58.72 39.40 37.30 69.56 57.78 6.78 6.14
Qwen2.5-7B-Instruct 49.65 44.98 34.67 32.16 64.22 48.56 6.70 4.77
Qwen3-8B 28.53 28.53 31.76 30.32 22.89 22.89 5.88 3.96
Atlas-Chat-9B 55.17 58.84 33.71 44.34 70.33 74.11 5.24 4.84
gemma-3-12b-it 61.17 60.00 38.59 35.66 75.78 64.89 8.76 7.09
AceGPT-13B-chat 45.45 40.68 35.06 32.40 38.78 36.44 6.10 4.83
jais-13b-chat 49.79 48.10 39.02 36.56 64.22 53.78 5.66 4.80
Table 1: Zero-shot performance of models on understanding and cultural knowledge evaluations. Metrics are accuracy for MMLU, HellaSwag, and Belebele, and a 0-10 correctness score for Palm. Bold values indicate the highest score among models comparable in size to ours (< 7B). Underlined values represent the highest score in the entire column, including larger models.
Understanding.
As Table 1 shows, NileChat demonstrates SoTA performance on the MMLU, HellaSwag, and Belebele benchmarks for both EGY and MOR when compared to similar size models. Specifically, NileChat surpasses its baseline model, Qwen2.5-3B-instruct, by a significant margin of $\sim$10 points across the majority of these tasks. Notably, NileChat also outperforms larger Arabic-focused models such as AceGPT-13B and Jais-13B. Furthermore, it achieves on-par performance with recent leading Arabic LLMs like ALLaM-7B, with a performance gap of less than 3 points on most tasks, and even surpasses it on certain benchmarks, including Belebele. Results for 3-shot are presented in Table D.1 and they show a similar trend to the zero-shot ones.
Cultural Knowledge.
As shown in Table 1, our approach significantly enhances cultural knowledge (Palm), enabling NileChat to achieve scores of 5.72 (EGY) and 5.86 (MOR), compared to baseline Qwen2.5-3B-instruct scores of 2.86 and 2.31, respectively. Among similarly sized models, ours achieves the highest performance on MOR and ranks second only to Gemma-3-4B for EGY. Although larger models such as Gemma-3-12B exhibit superior overall scores (EGY: 8.71, MOR: 7.09), NileChat notably surpasses AceGPT-7B and -13B on Moroccan cultural knowledge, despite their despite their claimed alignment with Arabic cultures. Additionally, it outperforms Atlas-chat-2B and -9B, models specifically fine-tuned for Moroccan dialects. These results support our claim that linguistic fluency alone—gained through supervised fine-tuning or pre-training on potentially biased, translated datasets—is insufficient for genuine cultural alignment with local communities.
Translation.
Table 2 summarizes the spBLEU scores from our zero-shot translation. Overall, NileChat achieves the highest average translation quality (spBLEU: 21.32), outperforming all evaluated models, including larger alternatives such as ALLaM-7B (20.60) and NLLB-200-3.3B (18.29). Specifically, on the Flores benchmark, NileChat demonstrates comparable performance to the similarly-sized NLLB-200-3.3B, with only a marginal 1-point spBLEU difference aggregated across MOR and EGY. Notably, NileChat surpasses even larger competitors in all translation directions, except when translating into MOR, where its performance matches that of Atlas-Chat-9B—a larger, single-dialect-focused model that is 3X larger.
On our in-house, human-curated dataset—which closely represents authentic speech patterns from local populations—NileChat significantly outperforms all baselines, including NLLB-200-3.3B, in all translation directions for both EGY and MOR. This real-world evaluation emphasizes the effectiveness of our strategy to incorporate local linguistic and cultural elements into synthetic data generation, enriching the pre-training data with diverse dialectal expressions and vocabulary. Detailed results for both zero-shot and 4-shot translation experiments are provided in Table D.2.
Model Flores-200 In-House Data Average
XX $\rightarrow$ $\rightarrow$ XX XX $\rightarrow$ $\rightarrow$ XX
EGY MOR EGY MOR EGY MOR EGY MOR
Less than 7B Qwen3-1.7B 14.75 10.89 19.51 15.47 11.41 4.36 15.63 6.32 12.29
ar-stablelm-2-chat 14.35 7.07 11.10 9.72 9.23 2.92 11.23 7.73 9.17
Atlas-Chat-2B 15.20 13.40 21.39 21.11 5.36 7.83 14.52 13.54 14.05
Llama-3.2-3B-Instruct 14.25 9.15 19.28 15.54 10.67 3.16 13.61 4.87 11.32
gemma-3-4b-it 9.27 5.22 12.46 10.13 3.01 0.60 16.89 5.25 7.86
Qwen3-4B 17.93 11.64 20.03 18.90 13.09 4.44 20.72 8.52 14.41
NLLB-200-3.3B 23.93 15.37 25.84 26.57 16.77 7.49 18.90 11.43 18.29
Qwen2.5-3B-Instruct 15.14 11.27 20.52 17.37 9.91 4.19 19.24 7.83 13.18
NileChat (3B) 23.60 16.41 25.74 25.56 22.02 12.34 26.50 18.39 21.32
More than 7B AceGPT-7B-chat 18.02 11.33 21.11 17.46 14.73 4.95 20.10 7.47 14.40
ALLaM-7B-Instruct 23.91 15.88 24.74 23.19 19.98 9.16 29.40 18.51 20.60
Qwen2.5-7B-Instruct 14.41 10.23 19.81 18.95 10.43 4.10 20.92 8.80 13.46
Qwen3-8B 20.03 13.86 22.56 21.33 13.38 4.73 24.14 9.27 16.16
Atlas-Chat-9B 18.20 16.89 24.92 26.29 5.36 7.68 17.35 15.23 16.49
gemma-3-12b-it 13.01 4.89 19.05 19.54 7.86 2.45 24.51 12.38 12.96
AceGPT-13B-chat 19.48 14.02 22.81 19.84 15.54 5.56 23.51 9.52 16.29
jais-13b-chat 8.80 4.29 15.77 17.12 10.83 4.02 19.19 12.47 11.56
Table 2: Zero-shot translation performance (spBLEU) on the Flores and in-house datasets. XX $\rightarrow$ EGY and XX $\rightarrow$ MOR denote average over target languages EGY and MOR, respectively. Conversely, EGY → XX and MOR → XX indicate average over EGY and MOR as source languages. Bold values highlight the top score among models with fewer than 7 billion parameters. Underlined values indicate the highest score overall in each column. Detailed results are in Table D.2.
Figure 3: Average SVA scores of evaluated models across societal value dimensions for Egypt and Morocco.
Value Alignment.
Figure 3 illustrates the results of value alignment evaluation based on the WVS. NileChat demonstrates substantial improvements over the baseline across most societal-value dimensions for both Moroccan and Egyptian contexts. Specifically, for Morocco, NileChat surpasses the baseline in all dimensions except Religious Values and the Index of Postmaterialism. Similarly, for Egypt, it outperforms the baseline across all dimensions except Political Interest and Political Participation, and the Index of Postmaterialism. These findings indicate that our approach—where a teacher LLM engages in role-playing by generating diverse text genres through personas embodying local community values—successfully steers responses towards culturally aligned positions. In a broader comparative analysis against all evaluated models, ours achieves the best results for Morocco across several dimensions and remains competitive in others. For Egypt, NileChat notably excels in Perceptions of Migration, Political Culture and Political Regimes, Happiness and Wellbeing, and Perceptions about Science and Technology, though models such as Jais-13B and ALLaM-7B show slightly stronger performance in certain other dimensions.
How many pre-training tokens are needed to reach good performance for a new language?
Figure 4 shows the performance evolution of NileChat during the pre-training phase on Belebele and translation tasks. The charts show that the model starts to get a large boost in these tasks during the first 10B tokens and then continues to slightly increase until it becomes steady after around 60B tokens.
Figure 4: Evolution of model performance during pre-training, measured by the number of tokens processed.
5 Conclusion
We introduced a novel methodology for adapting LLMs to specific communities by deeply integrating their unique linguistic characteristics, cultural heritage, and societal values. Our approach leverages a teacher model proficient in generating low-resource languages to enable: (i) translation for the incorporation of community-specific language and (ii) controlled generation and retrieval mechanisms for the authentic inclusion of cultural heritage and values. We validated our methodology using the Moroccan and Egyptian Arabic dialects as testbeds by developing NileChat, an LLM covering these two dialects. Comprehensive evaluations on understanding, translation, and cultural alignment benchmarks demonstrate that our method significantly enhances the baseline LLM’s performance in capturing target language nuances and cultural values. Notably, NileChat also outperform existing Arabic-aware LLMs. Our method offers a promising research direction for fostering inclusivity of diverse local communities within LLM development, thereby emphasizing the critical role of such an inclusion in the broader democratization of this technology.
Limitations
  • Teacher Model Dependency for Low-Resource Languages: Our method’s reliance on a teacher model proficient in generating even low-resource target languages may not hold for extremely under-resourced languages (e.g., Berber, Malayo-Polynesian varieties) Team et al. (2022), potentially limiting its applicability in such contexts.
  • Supervised Fine-Tuning Data: SFT phase predominantly utilized translated data due to resource constraints. This reliance on translated, rather than native, data for SFT might impact the model’s nuanced performance in the target languages and their corresponding cultures.
  • Susceptibility to Hallucination: As a 3B parameter model, our LLM is relatively small, rendering it more prone to hallucination and the generation of inaccurate or incomplete information compared to larger architectures Wei et al. (2022).
  • Computational Cost of Synthetic Data Generation: The process of generating synthetic data is computationally intensive, particularly when employing large teacher models (e.g., Command-R+, a 104B parameter model requiring substantial GPU resources: 4x80GB). This challenge is amplified by the autoregressive generation of long documents from extensive input contexts (e.g., articles, persona descriptions, cultural concepts) restricting the scale of this approach for more languages.
  • Absence of Explicit Safety Alignment: The model has not undergone dedicated safety alignment. While trained on curated datasets (Wikipedia, educational, news) largely devoid of toxic content and leveraging a safety-aligned teacher LLM, specific safety tuning is acknowledged as important future work.
  • Limited generation of subtle details. While the controlled generation uses multiple sources (WVS, news, Wikipedia, TV scripts), the generated texts are limited in terms of the very subtle cultural nuances, implicit knowledge, humor, or sarcasm that are often not explicitly stated in these source materials (wikipedia and news articles).
Ethics Statement
Our work contributes to the development of inclusive, linguistically, and culturally diverse LLMs capable of serving varied communities. While we generate our pre-training and instruction-tuning data using a teacher LLM, this process is critically informed by ground-truth cultural values survey data from the communities of interest and local context to control the generation. This approach aims to imbue our models with specific cultural nuances relevant to these communities.
As our evaluations demonstrate, the resulting models exhibit reasonable alignment with the cultural heritage and values of our target communities and can produce fluent text in their respective dialects. Despite these advancements, we have not conducted explicit safety alignment procedures for these models. Consequently, we strongly recommend thorough testing and further safety evaluations before any deployment in real-world scenarios.
Acknowledgments
We acknowledge support from Canada Research Chairs (CRC), the Natural Sciences and Engineering Research Council of Canada (NSERC; RGPIN-2018-04267), the Social Sciences and Humanities Research Council of Canada (SSHRC; 895-2020-1004; 895-2021-1008), Canadian Foundation for Innovation (CFI; 37771), Digital Research Alliance of Canada, [^6] https://alliancecan.ca and UBC Advanced Research Computing-Sockeye. [^7] https://arc.ubc.ca/ubc-arc-sockeye
Appendix A Data
Language Data category Data source Nature # of words
Dialectal data Egypt General Wikipedia Real 128.71M
MT fineweb-EDU Synthetic 2.08B
LHV Synthetic 398.89M
Fineweb2 Real 430.46M
Arabizi MT fineweb-EDU (Arabizi) & LHV Synthetic 206.49M
Morocco General Wikipedia Real 1.67M
Translated fineweb-EDU Synthetic 2.02B
LHV Synthetic 207.41M
Fineweb2 Real 1.64B
Arabizi MT fineweb-EDU (Arabizi) & LHV Synthetic 467.30M
MSA Egypt Cultural Brave API Real 74.67M
General Local News Real 346.79M
Morocco Cultural Brave API Real 23.08M
General Local News Real 220.16M
General General Fineweb2 Real 28.80B
Wikipedia Real 318.62M
English General General Fineweb-EDU Real 51.57B
French General General Fineweb2 Real 9.42B
Code & Math Code & Math
MathGenie/MathCode-Pile
macrocosm-os/code-parrot-github-code
Real 818.35M
Table A.1: Distribution of the final pre-training data mixture by language, nature (synthetic vs. real), and word count per dataset. Bold rows highlight data generated via our proposed augmentation pipeline.
Dataset name Language Source # of Instructions
Darija-SFT-Mixture MOR (Arabic) Atlas-Chat 458,155
TÜLU-V2-mix EGY (Arabic) Ours (MT) 178,109
SmolTalk MOR (Arabic) Ours (MT) 192,266
MOR (Arabizi) Ours (MT) 93,419
EGY (Arabic) Ours (MT) 195,260
EGY (Arabizi) Ours (MT) 93,181
French Ours (MT) 99,468
MSA Ours (MT) 96,933
English SmolTalk 149,124
ORCA MSA + dialects Ours (Converted) 460,203
Dolphin Ours (Converted) 425,703
Cultural instructions MOR (Arabic) Ours (Synthetic) 25,159
EGY (Arabic) Ours (Synthetic) 107,428
Table A.2: Distribution of the final instruction and response data mixture by language and number of instructions per dataset. ’Ours’ refers to datasets we created via machine translation (MT) or by converting existing datasets into an instruction/response format.
Table A.3: Examples of culturally aware and dialectally diverse texts from various genres, generated by our teacher model. The model was provided with input representing a specific persona, local context, and a local linguistic cue following the methodology described in Section 3.1.
Appendix B Training details
B.1 Pre-training
Motivations for choosing Qwen-2.5-3B as our backbone model. We select Qwen-2.5-3B as our base model to continue pretrain for two primary reasons: its competitive performance on MSA tasks and good tokenizer compression ratio on Arabic dialect texts. At the time of the selection, the Gemma 2 Team et al. (2024) and Qwen 2.5 base models showed the best performance in MSA. Also, both of their tokenizers have a good compression ratio for Arabic text in both standard and dialectal forms. Our analysis shows a ratio between 2.7 and 2.8 for Gemma, while Qwen 2.5 has a compression ratio between 2.9 and 3.1. Our final choice of Qwen 2.5 was based on its better performance on MSA.
Details for continued model pretraining. We continue the pre-training of Qwen-2.5-3B on our curated pre-training dataset. Subsequently, the model is fully fine-tuned for one epoch using a sequence length of 4,096. To optimize the learning process, the learning rate is linearly decayed from $5\times 10^{-6}$ to $5\times 10^{-7}$. To mitigate overfitting, we apply a weight decay of $0.1$, and gradient norms are clipped at a maximum value of $1.0$. The training is performed on a cluster of 4$\times$×A100 80GB GPUs.
B.2 Supervised fine-tuning
To enhance model robustness and facilitate effective merging, we augmented each dialect-specific dataset with a shared multilingual corpus, comprising English SmolTalk, MSA SmolTalk, French SmolTalk, and additional data from the ORCA and Dolphin datasets. Each dialect-specific model was trained for two epochs with a sequence length of 4,096 tokens, using a learning rate that linearly decayed from $7\times 10^{-6}$ to $7\times 10^{-7}$.
Table B.1 compares the SFT model with models fine-tuned on Egyptian and Moroccan datasets individually, as well as with our final merged model, NileChat. NileChat performs well on tasks for both EGY and MOR. The MOR-specific model also demonstrates strong performance on both MOR and, to some extent, EGY tasks. In contrast, the EGY-specific model does not perform well on MOR tasks.
Table B.1 compares the SFT model with models fine-tuned on Egyptian and Moroccan datasets individually, as well as with our final merged model, NileChat. NileChat performs well on tasks for both EGY and MOR. The MOR-specific model also demonstrates strong performance on both MOR and, to some extent, EGY tasks. In contrast, the EGY-specific model does not perform well on MOR tasks. We relate this observed asymmetry to the linguistic characteristics of the dialects relative to MSA. During the SFT phase, each dialect-specific dataset was augmented with a shared multilingual corpus which included MSA data (e.g., MSA SmolTalk, and data from the ORCA and Dolphin datasets). It is plausible that EGY is linguistically closer to MSA compared to the MOR, which is often considered more distant from MSA due to influences such as Berber and French. Consequently, the MOR-tuned model, having been exposed to this shared MSA data, might more effectively leverage this MSA knowledge to generalize to EGY tasks. Conversely, the greater linguistic divergence of the Moroccan dialect from MSA could make it more challenging for the EGY-tuned model to transfer its learning, including the MSA component, to the distinct features of the Moroccan dialect.
Belebele Flores In-house
[…]
NileChat-EGY 64.44 70.89 43.85 23.10 11.93 36.93
NileChat-MOR 70.67 72.56 39.94 37.45 30.82 29.98
NileChat 70.33 72.67 44.37 33.89 28.67 37.52
Table B.1: Comparison of the performance of the Egyptian SFT model (NileChat-EGY), the Moroccan SFT model (NileChat-MOR), and their merged version, NileChat, on Belebele (accuracy), Flores (chrf), and In-house parallel data (chrf).
Appendix C Evaluation Setup
C.1 Evaluation Tasks
Full list of the 13 categories of WVS questions. Economic Values (EcoV); Ethical Values (EthV); Happiness and Wellbeing (HW); Index of Postmaterialism (IP); Perceptions about Science and Technology (PST); Perceptions of Corruption (PC); Perceptions of Migration (PM); Perceptions of Security (PS); Political Culture and Political Regimes (PCPR); Political Interest and Political Participation (PIPP); Religious Values (RV); Social Capital, Trust, and Organizational Membership (SCTOM); and Social Values, Norms, and Stereotypes (SVNS).
C.2 Baselines
We evaluate our model NileChat against a set of 17 LLMs that are Arabic-aware; some of these 17 models are also aligned to Arabic dialects. These models are from the following model families: ALLaM Bari et al. (2025), Jais Sengupta et al. (2023), Atlas-Chat Shang et al. (2025), ar-stablelm-2-chat Alyafeai et al. (2024), Gemma-3 Team et al. (2025b), Qwen-2.5 Qwen et al. (2025), Qwen3 Yang et al. (2025) (non-thinking mode), and Llama-3.2 Grattafiori et al. (2024). The full list of models, including their corresponding size and release date, are presented in Table C.1.
Model Name Size Release Date
Less than 7B
Qwen3-1.7B 1.7 Apr. 2025
ar-stablelm-2-chat 1.6 Jul. 2024
Atlas-Chat-2B 2.6 Sep. 2024
Llama-3.2-3B-Instruct 3.2 Sep. 2024
gemma-3-4b-it 4.3 Mar. 2025
Qwen3-4B 4 Apr. 2025
NLLB-200-3.3B 3.3 Jul. 2022
Qwen2.5-3B-Instruct 3.1 Sep. 2024
More than 7B
AceGPT-7B-chat 7 Dec. 2023
ALLaM-7B-Instruct 7 Feb. 2025
Qwen2.5-7B-Instruct 7.6 Apr. 2025
Qwen3-8B 8.2 Apr. 2025
Atlas-Chat-9B 9.2 Sep. 2024
gemma-3-12b-it 12.2 Mar. 2025
AceGPT-13B-chat 13 Dec. 2023
jais-13b-chat 13 Aug. 2023
Table C.1: The LLMs used for comparison against NileChat in this evaluation were selected from a list of Arabic-aware models. Each LLM is listed with its corresponding size (in billion parameters) and release date. We utilized the instruct version for all LLMs except for NLLB, which is a machine translation-specific model.
Appendix D Full results
Model MMLU HellaSwag Belebele
EGY MOR EGY MOR EGY MOR
Less than 7B Qwen3-1.7B 28.53 28.53 28.07 27.33 22.89 22.89
ar-stablelm-2-chat 39.54 38.32 34.33 33.40 24.22 22.78
Atlas-Chat-2B 42.65 45.06 29.62 34.78 54.67 59.00
Llama-3.2-3B-Instruct 31.10 30.92 28.86 28.39 49.67 40.89
gemma-3-4b-it 46.32 46.60 34.26 32.53 61.44 52.11
Qwen3-4B 28.59 28.52 30.21 29.53 22.89 22.89
Qwen2.5-3B-Instruct 35.71 37.67 31.17 29.62 61.11 44.89
NileChat (3B) 58.20 58.62 38.29 40.35 78.11 73.78
More than 7B AceGPT-7B-chat 40.76 37.98 33.04 31.04 38.00 33.00
ALLaM-7B-Instruct 60.18 59.61 40.20 38.14 76.11 66.00
Qwen2.5-7B-Instruct 57.70 53.51 33.79 32.28 76.67 59.44
Qwen3-8B 28.53 28.53 31.72 30.95 22.89 22.89
Atlas-Chat-9B 57.17 60.27 34.75 44.47 78.44 79.33
gemma-3-12b-it 59.29 56.16 40.16 37.60 80.78 73.11
AceGPT-13B-chat 46.48 43.65 35.15 33.21 46.33 41.11
jais-13b-chat 49.33 48.28 38.99 37.45 59.89 53.78
Table D.1: 3-shot performance (accuracy) of models on understanding (MMLU, HellaSwag, and Belebele). Bold values indicate the highest score among models comparable in size to ours (<7B parameters). Underlined values represent the highest score in the entire column, including larger models. Results for zero-shot are presented in Table 1, Section 4.2.
Flores Dataset Performance
Model MOR $\rightarrow$ EGY MOR $\rightarrow$ ENG MOR $\rightarrow$ FRA EGY $\rightarrow$ MOR EGY $\rightarrow$ ENG ENG $\rightarrow$ EGY ENG $\rightarrow$ EGY FRA $\rightarrow$ EGY Average
sB0 sB4 cF0 cF4 sB0 sB4 cF0 cF4 sB0 sB4 cF0 cF4 sB0 sB4 cF0 cF4 sB0 sB4 cF0 cF4 sB0 sB4 cF0 cF4 sB0 sB4 cF0 cF4 sB0 sB4 cF0 cF4 sB0 sB4 cF0 cF4
Less than 7B Qwen3-1.7B 19.49 19.80 36.32 36.45 14.76 15.45 40.21 40.79 12.17 12.90 35.08 35.83 19.38 19.31 36.23 36.22 19.65 20.48 46.49 47.05 7.32 7.77 22.02 23.42 10.02 10.62 25.98 27.55 5.98 6.63 19.60 21.69 13.59 14.12 32.74 33.62
ar-stablelm-2-chat 18.29 18.57 34.31 34.46 7.36 7.03 28.97 28.47 3.51 2.85 19.36 18.08 14.16 16.35 29.25 33.32 8.05 7.93 31.12 30.67 3.25 11.51 8.22 29.25 10.41 18.68 22.89 35.77 3.81 9.55 9.06 26.13 8.60 11.56 22.90 29.52
Atlas-Chat-2B 19.67 20.03 36.67 36.90 24.64 26.18 49.36 50.32 19.03 20.18 41.53 42.49 19.35 19.88 36.88 37.65 23.44 25.41 49.17 50.28 12.06 14.89 29.66 33.48 10.74 15.16 27.45 33.04 8.80 11.72 25.73 29.77 17.22 19.18 37.06 39.24
Llama-3.2-3B-Instruct 18.75 10.79 35.15 21.71 15.71 0.08 38.75 0.85 12.16 0.54 29.55 3.40 17.67 14.00 34.59 28.64 20.89 0.06 45.19 0.64 5.86 0.91 16.92 5.21 9.74 0.77 24.54 5.29 3.92 1.12 15.14 5.64 13.09 3.53 29.98 8.92
gemma-3-4b-it 16.89 21.11 33.00 37.78 7.99 24.48 17.81 49.50 5.52 23.80 11.28 46.04 12.86 16.31 29.10 33.91 12.05 29.04 24.62 54.24 1.88 13.05 4.85 31.03 1.65 20.13 3.24 37.60 0.93 10.56 3.01 28.17 7.47 19.81 15.86 39.78
Qwen3-4B 20.41 20.57 37.17 37.15 18.98 20.67 44.68 46.07 17.30 18.39 39.96 41.33 16.06 18.86 31.90 35.84 23.99 25.75 50.47 51.83 11.43 11.30 27.93 27.98 15.46 15.83 33.00 33.20 7.44 9.54 21.46 25.90 16.38 17.62 35.82 37.41
NLLB-200-3.3B 20.92 - 38.74 - 30.89 - 53.64 - 27.89 - 48.69 - 17.06 - 35.14 - 34.62 - 58.07 - 17.46 - 34.89 - 26.93 - 43.86 - 11.58 - 28.90 - 23.42 - 42.74 -
Qwen2.5-3B-Instruct 18.51 19.71 35.20 36.32 18.61 19.83 44.22 44.99 15.00 16.22 38.56 39.45 16.98 17.90 34.15 34.91 24.07 25.08 50.45 50.99 9.36 8.48 25.17 25.14 11.78 11.43 28.05 28.44 7.47 7.18 22.79 23.32 15.22 15.73 34.83 35.45
NileChat 23.81 22.68 40.25 39.25 28.81 29.50 52.98 53.43 24.05 23.49 46.30 45.56 20.90 20.51 38.22 37.64 30.58 31.90 54.53 56.19 15.15 17.90 30.76 36.00 23.39 25.03 40.37 41.94 13.18 14.91 29.61 32.52 22.49 23.24 41.63 42.82
More than 7B AceGPT-7B-chat 19.92 18.93 36.91 35.72 17.67 20.04 44.09 45.71 14.78 15.92 38.08 38.87 19.11 17.99 36.04 35.24 23.11 25.81 50.28 51.96 8.93 12.29 21.59 29.72 16.11 18.25 33.32 35.56 5.95 10.28 16.65 27.07 15.70 17.44 34.62 37.48
ALLaM-7B-Instruct 24.63 25.12 40.93 41.39 26.02 28.13 51.65 52.92 18.91 20.80 42.64 44.04 19.85 19.53 37.32 37.00 29.63 31.82 55.53 56.75 14.63 16.56 32.38 34.79 23.18 25.93 39.85 42.70 13.17 14.09 30.75 32.15 21.25 22.75 41.38 42.72
Qwen2.5-7B-Instruct 16.28 18.87 32.19 35.08 21.59 23.62 47.28 48.55 18.98 19.96 42.12 42.93 13.24 16.09 29.34 32.91 26.39 27.83 52.72 53.74 9.73 11.19 25.32 28.28 12.54 15.00 27.80 31.34 7.73 9.75 22.71 26.46 15.81 17.79 34.93 37.41
Qwen3-8B 21.10 21.41 37.84 38.13 22.41 24.33 48.04 49.51 20.46 21.91 43.17 44.32 17.82 18.35 34.63 35.40 27.30 29.36 53.32 54.77 13.60 12.56 31.05 29.63 18.95 18.83 36.79 36.72 10.16 10.27 26.33 26.86 18.98 19.63 38.90 39.42
Atlas-Chat-9B 19.44 20.52 36.32 37.29 31.29 32.75 54.19 55.61 28.14 28.82 49.35 49.85 18.86 20.36 36.78 37.91 30.98 32.83 55.04 56.07 17.29 18.91 35.83 37.42 16.95 19.69 34.35 37.25 14.53 15.59 32.76 33.92 22.19 23.68 41.83 43.16
gemma-3-12b-it 19.67 24.48 36.28 41.24 23.68 30.29 47.75 54.47 15.28 29.99 26.64 51.36 10.73 17.30 24.98 36.28 27.38 32.71 53.46 57.42 2.51 14.60 6.56 34.00 6.34 24.27 11.17 41.69 1.45 12.47 3.70 31.30 13.38 23.26 26.32 43.47
AceGPT-13B-chat 20.14 19.58 37.04 36.57 20.82 24.30 46.12 49.11 18.58 20.52 40.45 42.77 19.54 17.99 36.62 35.50 26.08 29.14 52.52 54.71 13.02 13.42 29.23 31.52 18.81 20.18 36.04 37.93 9.51 11.72 23.19 29.39 18.31 19.61 37.65 39.69
jais-13b-chat 13.50 20.28 26.36 37.46 19.73 27.40 37.81 51.80 18.14 18.40 39.16 40.95 8.90 13.42 22.22 31.25 22.63 30.07 41.73 54.56 2.46 8.55 6.08 25.73 4.09 18.77 9.56 36.94 1.50 7.26 3.25 23.70 11.37 18.02 23.27 37.80
In-House Dataset Performance
Model EGY $\rightarrow$ ENG EGY $\rightarrow$ MSA ENG $\rightarrow$ EGY ENG $\rightarrow$ MOR MOR $\rightarrow$ EGY MOR $\rightarrow$ MSA MSA $\rightarrow$ EGY MSA $\rightarrow$ MOR Average
sB0 sB4 cF0 cF4 sB0 sB4 cF0 cF4 sB0 sB4 cF0 cF4 sB0 sB4 cF0 cF4 sB0 sB4 cF0 cF4 sB0 sB4 cF0 cF4 sB0 sB4 cF0 cF4 sB0 sB4 cF0 cF4 sB0 sB4 cF0 cF4
Less than 7B Qwen3-1.7B 12.14 13.14 31.46 31.98 19.11 20.32 34.24 35.09 4.63 4.96 15.10 15.43 2.52 2.64 11.14 10.99 6.19 6.12 22.58 23.01 6.45 6.04 19.25 18.12 18.20 18.26 34.20 33.86 6.20 6.07 18.38 17.94 9.43 9.69 23.29 23.30
ar-stablelm-2-chat 7.89 6.74 21.92 20.28 14.56 13.71 27.71 24.23 3.57 10.99 7.50 22.25 0.65 3.69 1.45 14.61 6.95 6.15 20.63 19.05 8.50 7.28 21.32 18.59 14.89 14.28 28.14 27.20 5.19 3.49 15.82 13.12 7.77 8.29 18.06 19.92
Atlas-Chat-2B 13.52 16.21 32.04 34.25 15.51 18.19 30.39 32.34 3.67 5.15 15.29 17.15 6.81 8.40 21.76 23.79 14.96 15.85 34.83 35.82 12.12 12.76 28.38 28.48 7.05 11.91 20.72 26.75 8.86 10.11 24.35 25.20 10.31 12.32 25.97 27.97
Llama-3.2-3B-Instruct 12.97 0.38 28.40 1.37 14.24 4.80 25.90 8.81 5.60 2.41 16.44 8.16 2.03 0.98 10.14 5.63 4.74 0.23 17.03 1.01 5.00 1.98 14.55 6.79 15.73 7.97 31.80 18.17 4.30 2.46 14.68 11.25 8.08 2.65 19.87 7.65
gemma-3-4b-it 16.24 22.91 32.96 43.72 17.53 26.98 30.61 42.52 0.52 13.66 1.61 28.41 0.49 5.67 1.90 18.59 6.10 12.65 16.81 33.20 4.41 12.67 11.22 28.30 5.50 19.30 11.35 35.56 0.72 6.16 2.13 18.92 6.44 15.00 13.57 31.15
Qwen3-4B 19.21 18.42 39.11 38.95 22.22 23.63 37.84 38.93 8.19 8.16 20.04 20.48 3.41 3.40 13.35 13.57 9.18 8.54 26.74 26.42 7.86 8.51 21.12 21.64 17.98 19.72 33.83 35.71 5.47 6.15 16.43 18.51 11.69 12.07 26.06 26.78
NLLB-200-3.3B 21.48 - 41.14 - 16.33 - 34.36 - 14.20 - 29.78 - 6.28 - 19.86 - 14.12 - 32.44 - 8.75 - 26.11 - 19.34 - 36.68 - 8.70 - 23.10 - 13.65 - 30.43 -
Qwen2.5-3B-Instruct 19.44 19.38 39.84 39.59 19.04 20.48 34.07 34.75 6.34 6.58 17.94 18.97 3.38 3.65 13.01 14.36 8.15 9.49 26.02 27.75 7.52 7.68 20.87 20.82 13.47 16.82 28.91 33.21 5.00 5.40 17.04 17.88 10.29 11.19 24.71 25.92
NileChat 26.24 27.49 46.46 48.10 26.77 31.65 42.52 46.40 17.67 18.92 33.37 34.87 10.74 11.72 25.07 26.84 19.22 20.47 39.83 41.46 17.55 19.28 34.23 36.55 26.37 29.54 42.90 46.68 13.94 15.41 29.37 30.75 19.81 21.81 36.72 38.95
More than 7B AceGPT-7B-chat 19.27 21.44 39.36 40.93 20.94 21.99 35.99 37.51 9.24 10.57 22.39 25.25 3.36 3.69 12.76 15.12 7.80 9.95 26.08 27.97 7.15 8.23 20.20 22.17 20.22 20.47 37.14 37.05 6.53 6.00 19.31 18.98 11.81 12.79 26.65 28.12
ALLaM-7B-Instruct 26.31 28.98 48.32 49.57 32.49 32.85 48.22 47.93 16.69 17.78 32.85 34.12 8.05 9.32 22.65 24.47 18.61 19.97 39.44 40.36 18.41 19.60 34.86 36.02 23.28 26.15 41.09 43.79 10.28 11.13 25.95 26.49 19.26 20.72 36.67 37.84
Qwen2.5-7B-Instruct 22.70 21.40 43.05 42.34 19.14 22.90 33.08 36.71 7.65 7.91 19.04 20.16 3.65 4.50 13.87 16.04 9.87 10.55 28.38 29.17 7.72 9.05 20.84 22.79 13.21 16.00 28.10 31.91 4.55 5.74 15.67 18.44 11.06 12.26 25.25 27.20
Qwen3-8B 22.21 23.62 42.36 44.34 26.06 27.35 41.28 42.38 9.04 9.52 22.25 23.18 3.67 4.50 14.64 15.68 10.23 11.03 28.90 29.63 8.30 9.16 21.79 23.18 17.71 19.49 33.99 35.90 5.79 6.12 17.83 18.83 12.88 13.85 27.88 29.14
Atlas-Chat-9B 18.98 21.19 38.91 41.38 15.72 20.68 31.29 35.84 4.53 7.74 17.12 21.46 7.19 10.60 22.68 26.51 16.77 19.19 37.84 40.15 13.70 15.38 29.86 31.29 6.19 12.05 19.20 27.01 8.16 12.53 23.70 28.12 11.41 14.92 27.58 31.47
gemma-3-12b-it 23.12 28.51 46.31 48.58 25.91 32.60 40.06 48.67 6.59 17.70 14.15 34.10 1.83 7.95 6.29 23.15 12.49 18.35 33.40 39.90 12.26 17.54 23.93 34.51 9.14 26.48 18.35 43.51 3.06 10.73 8.83 26.07 11.80 19.98 23.91 37.31
AceGPT-13B-chat 22.36 25.37 42.53 45.47 24.67 27.42 40.60 42.11 10.76 12.56 25.54 27.49 4.27 5.17 15.26 17.90 10.56 12.24 28.12 31.28 8.47 10.30 22.28 24.78 20.32 22.83 37.61 39.75 6.86 7.31 20.34 20.72 13.53 15.40 29.04 31.19
jais-13b-chat 16.71 24.94 33.16 43.30 21.68 28.15 36.23 43.52 6.45 14.19 14.78 30.35 2.47 6.35 7.42 20.10 12.06 18.56 26.34 38.44 12.89 18.74 28.36 34.79 15.20 22.74 30.81 39.83 5.57 8.83 16.16 23.75 11.63 17.81 24.16 34.26
Table D.2: Detailed translation evaluation results on Flores and In-House datasets. Metrics: spBLEU (sB) and chrF (cF) for 0-shot (0) and 4-shot (4) scenarios. Bold values highlight the top score among models with fewer than 7 billion parameters. Underlined values indicate the highest score overall in each column.
Appendix E Prompts
The provided figures showcase diverse prompts for language models targeting low-resource languages. Figure 5 translates English educational content into conversational dialectal Arabic, while Figure 6 converts dialectal Arabic script to Arabizi. English instructions are translated to dialectal Arabic using the prompt in Figure 7. For content generation, Figure 8 guides the model to create culturally relevant dialectal Arabic text based on a given persona and context. Figure 9 focuses on summarizing detailed persona descriptions concisely. Finally, Figure 10 instructs an LLM to generate practical dialectal Arabic question-answer pairs in JSON format from provided text.
Translate the following text from English to Egyptian Arabic. Ensure that all words are in Egyptian Arabic, and do not use any Modern Standard Arabic (MSA). Keep the translation casual, conversational, and reflective of how Egyptians would naturally speak in everyday situations. Avoid any formal or classical language structures. Translate only the input paragraph and dont add anything else in your output.
English: {English_text}
Figure 5: The translation prompt used with a teacher model to convert English educational pre-training data to a low-resource target language. The placeholder ’{English_text}’ represents the input English text.
Write the following Moroccan dialectal Arabic text in Moroccan Arabizi. Ensure that all words are written in Moroccan Arabizi. Keep the text casual, conversational, and reflective of how Moroccans would naturally write in everyday situations using Arabizi. Translate only the content keys in the following JSON, and output a json of the same format:
{JSON_OBJECT}
Figure 6: The prompt used with our teacher LLM to convert dialectal Arabic text written in Arabic script into Arabizi. The placeholder {JSON_OBJECT} represents the input text formatted as a JSON object.
Translate the following text from English to Moroccan Arabic. Ensure that all words are in Moroccan Arabic, and do not use any Modern Standard Arabic (MSA). Keep the translation casual, conversational, and reflective of how Moroccans would naturally speak in everyday situations. Avoid any formal or classical language structures. Translate only the content keys in the following JSON, and output a json of the same format:
{JSON_OBJECT}
Figure 7: The translation prompt used with a teacher model to convert SmolTalk and TULU instructions data to a low-resource target language. The placeholder ’{JSON_OBJECT}’ represents the input text.
Act as the following person: {persona_description} Act like you are {person_Name} and write a {text_genre} in Egyptian dialect, using colloquial Arabic script as spoken in Egypt and not Modern Standard Arabic (MSA). Use this context and use the information provided in it while writing the {text_genre}:
{context}
Make sure to follow these conditions:
1. Rely on the provided context when writing the {text_genre}.
2. Ensure that the written {text_genre} reflects the cultural background, values, and worldview of {person_Name}.
3. Dont write the personas description. I want you to focus only on the provided context when writing while reflecting the perosnas background.
Note: Ensure that all words are in Egyptian Arabic, and do not use any Modern Standard Arabic (MSA). Keep the translation casual, conversational, and reflective of how Egyptians would naturally speak in everyday situations.
Figure 8: Prompt for generating culturally and values-aware text genres in low-resource languages, given a local persona description and a local cultural concept. The placeholders {persona_description}, {text_genre}, and {context} represent the persona description, the intended text genre to generate, and the cultural concept text, respectively.
I have the following persona description, I want you to write it in a concise manner keeping all the information, the output should be plain text, make sure to include all values, morals, and culture of the persona:
{PERSONA_DESCRIPTION}
Figure 9: Prompt for generating concise persona descriptions with a LLM utilizing a comprehensive description of a specific persona extracted from the WVS. The placeholder ’{PERSONA_DESCRIPTION}’ represents the input persona description.
You are an expert in Moroccan culture and language, with native-level fluency in Moroccan Arabic dialect written in Arabic script. Your task is to transform detailed articles or information (like recipes, historical accounts, cultural traditions, etc.) into practical, useful instruction-response pairs in authentic Moroccan Arabic dialect expressed in Arabic script.
## Task Description:
1. I will provide you with raw text content (like cooking recipes, cultural traditions, historical information, etc.)
2. Transform this content into practical instruction-response pairs where:
- Instructions ask how to do something, how something works, or how to understand something
- Responses provide thorough, practical explanations in Moroccan dialect
3. Focus on making these interactions useful for someone wanting to learn practical information
4. Output the results in structured JSON format
## Instruction Format Guidelines:
- For procedural content (recipes, crafts, etc.): Create "how to" questions
- Example: {example_1}
- For historical/factual content: Create "why" questions
- Example: {example_2}
- For cultural practices: Create "what is" or "how do we celebrate" questions
- Example: {example_3}
## Response Format Guidelines:
- Make responses detailed and practical
- Include specific steps for procedural content
- Use authentic Moroccan vocabulary and expressions
- Structure responses in a logical order
- For recipes or procedures, list steps clearly in the order they should be performed
- Include tips and warnings where appropriate
- Keep the tone conversational and helpful, as if explaining to a friend
## Language Guidelines:
- Use authentic Moroccan Arabic dialect written in Arabic script (not transliteration)
- Include typical Moroccan cooking/cultural terminology and expressions
- Use language as it would naturally be spoken, not literary Arabic
- Include common Moroccan filler words and expressions where natural
- For cooking terms or specialized vocabulary, use the actual terms Moroccans use.
## Output Format:
The output should be valid JSON with the following structure:
“‘json
"instruction_response_pairs": [
"instruction": "[practical instruction in Moroccan dialect]",
"response": "[detailed practical response in Moroccan dialect]"
"instruction": "[another practical instruction]",
"response": "[another detailed response]"
## Instructions:
1. Generate at least 3-5 practical instruction-response pairs based on the content I provide.
2. Make each instruction focus on a practical aspect (how-to, ingredients, common mistakes, variations, etc.).
3. Ensure responses are detailed enough to be practically useful.
4. The resulting JSON must be properly formatted and valid.
5. For specialized terms or techniques, use the authentic Moroccan terms.
6. When dealing with recipes or procedures, make sure the steps are in logical order.
7. If the article contains cultural context or historical information, include some of this in the responses to provide depth.
8. If the article does not have any relevent information to the Moroccan culture then return an empty list.
Now, Ill provide you with content, and youll transform it into practical instruction-response pairs in Moroccan dialect as specified above.
{Text}
Figure 10: Prompt for generating synthetic cultural dialectal question-answer pairs using a teacher model given raw text describing a local cultural heritage concept. The placeholder {Text} represents the raw text that is used to generate question-answer pairs from it.
Appendix F Samples of texts generated from NileChat
Table F.1: Sample responses from NileChat to prompts in Egyptian and Moroccan dialects, covering general and local cultural knowledge. Samples with green background color represent samples with correct responses, samples with red background color represent samples with not accurate answers.